### v1

@@ Abstract
Construction workers exert intense physical effort and experience serious safety and health risks in hazardous working environments. Quiet stance and kneeling are among the most common postures performed by construction workers during their daily work. This paper analyzes lower-limb joint influence on neural balance control strategies using the frequency behavior of the intersection point of ground reaction forces. To evaluate the impact of elevation and wearable knee exoskeletons on postural balance and welding task performance, we design and integrate virtual- and mixed-reality (VR/MR) to simulate elevated environments and welding tasks. A linear quadratic regulator-controlled triple- and double-link inverted pendulum model is used for balance strategy quantification in quiet stance and kneeling, respectively. Extensive multi-subject experiments are conducted to evaluate the usability of occupational exoskeletons in destabilizing construction environments. The quantified balance strategies capture the significance of knee joint during balance control of quiet stance and kneeling gaits. Results show that center of pressure sway area reduced up to 62% in quiet stance and 39% in kneeling for subjects tested in high-elevation VR/MR worksites when provided knee exoskeleton assistance. The comprehensive balance and multitask evaluation methodology developed aims to reveal exoskeleton design considerations to mitigate the fall risk in construction.
@@
Construction workers commonly perform tasks that require prolonged quiet stance or kneeling gaits on high elevations. Worker balance can be undermined by chronic knee injuries, musculoskeletal disorders and destabilizing visual perturbations caused by occupational activities. Wearable knee exoskeletons have evolved as promising interventions to reduce knee joint stress across a variety of work gaits in construction. Emerging technologies such as virtual- and mixed-reality (VR/MR) provide a platform to study underlying balance strategies workers employ to complete tasks when confronted by dynamic environment. The VR/MR- generated immersive elevated welding work sites in this study are leveraged to examine the effects of threatening visual stimuli, wearable exoskeletons and construction tasks on worker balance and performance. Intersection point height frequency analysis is used to observe and quantify the chosen neural balance strategy during various testing scenarios. Emphasis is placed on exploring the often-neglected role of the knee joint to facilitate research on knee-based balance augmenting exoskeleton controllers that can support subtle gaits like quiet stance and kneeling. The experimental results provide insight into the efficacy of knee exoskeletons in improving construction worker stability and task performance. The results highlight the need for a holistic approach to exoskeleton design and evaluation to ensure that developed solutions can be safely and successfully integrated into the workplace environment.
@@ I Introduction
Construction workers commonly perform tasks that require prolonged quiet stance or kneeling gaits on high elevations or cluttered environments. These awkward gaits in construction increase the risk of work-related musculoskeletal disorders (WMSDs), such as knee pain or knee osteoarthritis [1]. In both stance and kneeling gaits, human workers are inherently unstable in an upright posture, requiring continuous engagement of muscles to regulate joint torques and maintain postural stability [2, 3]. Although the neural balance mechanism at quiet stance has been extensively studied, the underlying control strategy remains elusive [4]. Previous studies have investigated the ankle or hip strategies during quiet standing [5, 6], and most of them used the single-link inverted pendulum model to analyze human balance strategies. The importance of the knee and hip joints at quiet stance was demonstrated in [7, 8]. However, the knee joint contribution to maintaining an upright stance is often overlooked. Few research work was reported for postural control of kneeling gaits. In [9], a comparative analysis of quiet kneeling and stance postures was conducted with a focus on the role of visual feedback during kneeling. A single inverted pendulum model was used in [6] to study the influences of elevation on the postural balance of kneeling gait.
@@
All of the works above primarily use the motion of the body’s center of mass (COM) or center of pressure (COP) to quantify the human balance performance. Metrics such as power spectral density (PSD) of the COP and acceleration were used to compare and discriminate between quiet stance and kneeling [9, 6]. The metrics solely derived from COP or COM are limited in capturing the balance dynamics and control as they are unable to account for both translational and rotational body accelerations. To address these limitations, the study in [10, 4] introduced the intersection point (IP) of the ground reaction forces (GRF) during quiet stance. In [11], a double-inverted pendulum (DIP) model was leveraged to quantify the frequency-dependence of the IP height during quiet stance. The study focused on quantifying ankle and hip joint strategies and emphasized the importance of multiple joint-level contributions to overall balance control.
@@
Among the diverse environmental factors affecting balance control for construction workers, elevation is notably significant [12, 13, 14]. Acrophobia exacerbates this challenge, as it leads individuals to overestimate heights, intensifying fear and triggering physiological reactions that undermine balance [15]. Additionally, visual and environmental perturbations further compromise postural stability [16]. To address these challenges, wearable devices like knee exoskeletons have emerged as pivotal fall risk interventions [17, 18, 19, 20]. However, little work has been done on knee-based balance augmentation for quiet stance and kneeling. Virtual- and mixed-reality (VR/MR) has been used to simulate high-elevation conditions and study neural balance control among construction workers [21, 6]. The work in [22] showed safety reinforcement training on construction workers’ fall risks using VR scenes. The work in [23] conducted neurophysiological assessments for welding task performance at virtual height to study cognitive effects on balance. A comprehensive review of VR/MR applications in construction safety can be found in [24].
@@
The goal of this study is to investigate the influence of lower-limb joints, particularly the knee joint, on neural balance control in quiet stance and kneeling gaits. We also evaluate the impact of elevation and wearable knee exoskeletons on postural balance and welding task performance. Similar to previous studies (e.g., [25, 26, 11]), a linear quadratic regulator (LQR) is used as the human neural balance controller. The triple- and double-link inverted pendulum (TIP and DIP) models are built to capture the joint-level effects for quiet stance and kneeling gaits, respectively. The models are used to quantify the observed neural balance control strategy from frequency characteristics of the IP heights. The LQR allows for the explicit consideration of different parameters on balance performance, facilitating the interpretation of results in terms of physiological and biomechanical implications [11]. Human subject experiments are conducted using incorporating VR/MR environments that simulate realistic welding worksites at varying elevations to evaluate the influence of elevation on postural balance and task performance. The subjects are outfitted with a wearable knee exoskeleton to test the usability and functionality of the intervention. The combination of experimental data and model analysis helps the extraction of physiologically plausible insights into the balance strategy chosen by subjects. Both objective metrics and human subjective feedback confirm that the exoskeletons have the potential to improve both balance and task execution efficiency in construction.
@@
The main contributions of this work are twofold. First, the study introduces new modeling and analysis for neural balance control that accounts for knee joint strategy with the IP height as the performance metric for quiet stance and kneeling. To the authors’ best knowledge, no previously reported work elucidates the involvement of knee joint strategy in quiet stance or kneeling gaits. Second, this study introduces a novel exploration of how elevation and the use of wearable exoskeletons impact neural balance control during quiet stance and kneeling. The quantitative findings offer valuable insights for the development of balance-enhancing wearable robotic interventions. Compared with the previous conference presentation [27] that focused only on quiet stance, this work extends to kneeling gaits by introducing the DIP model to investigate the joint-level effects on postural balance control. Additionally, the VR/MR simulating welding tasks to assess worker performance is newly developed.
@@
The rest of the paper is organized as follows. We introduce the multi-link models and neural balance control in Section II. The experimental setup and protocols are detailed in Section III. Data analysis and simulation pipelines are explained in Section IV. Section V presents the experimental and simulation results, followed by discussions in Section VI. Finally, we give the concluding summary in Section VII.
@@ II Multi-Link Models and Balance Control
In this section, we first present the inverted pendulum models that describe quiet stance and kneeling, respectively. Next, the IP for human postural balance is discussed followed by presentation of the LQR-based neural controller.
@@ II-A Biomechanical Models for Stance and Kneeling
Fig. 1 illustrates the welder’s stance posture and schematic consisting of three interconnected segments. The upper segment, which includes the head, arms, and trunk, is connected to the middle segment, representing the thigh, and the lower segment, representing the shank. In this study, only motion in the sagittal plane is considered. Fig. 1 shows the welder’s working pose when kneeling on the ground with both legs alongside the schematic of the DIP model. In the case of quiet stance, the ankle joint is assumed to be the pin joint on the ground, whereas for kneeling, the knee joint is assumed to be the pin joint on the ground and feet are assumed to support no weights. The joint angles and joint torques at the ankle, knee, and hip are denoted as $\theta_{i}$ and $\tau_{i}$, $i=1,2,3$, respectively. The masses and mass moments of inertia about the centers of each link are denoted as $m_{i}$ and $I_{i}$, $i=1,2,3$, respectively. The distance between the mass center and the lower end of each link is denoted as $l_{c_{i}}$, $i=1,2,3$, respectively.
Defining ${\boldsymbol{\theta}}_{t}=[\theta_{1}\;\theta_{2}\;\theta_{3}]^{T}$ and ${\boldsymbol{\theta}}_{d}=[\theta_{2}\;\theta_{3}]^{T}$, the equations of motion for the TIP/DIP models are written as
@@
Defining ${\boldsymbol{x}}_{i}=[{\boldsymbol{\theta}}_{i}^{T}\;{\boldsymbol{\dot{\theta}}}_{i}^{T}]^{T}$, $i=t,d$, from (1), we obtain
@@
The position vectors of the COM are denoted as ${\boldsymbol{r}}^{M}_{i}=[x^{m}_{i}\,z^{m}_{i}]^{T}$, $i=t,d$, for stance and kneeling gaits, respectively. For stance posture, $x^{m}_{t}=-l_{c1}\mathop{\rm s}\nolimits_{\theta_{1}}-l_{c2}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}}-l_{c3}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}$ and $z^{m}_{t}=l_{c1}\mathop{\rm c}\nolimits_{\theta_{1}}+l_{c2}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}}+l_{c3}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}$, where notations $\mathop{\rm s}\nolimits_{\theta_{1}}=\sin\theta_{1}$, $\mathop{\rm c}\nolimits_{\theta_{1}}=\cos\theta_{1}$ for $\theta_{1}$, and $\mathop{\rm s}\nolimits_{\theta_{i}+\theta_{j}}=\sin(\theta_{i}+\theta_{j})$ for angles $\theta_{i}$ and $\theta_{j}$, $i\neq j$, $i,j=1,2,3$. For quiet stance, we define the Jacobian matrix of the COM to ${\boldsymbol{x}}_{i}$ as ${\boldsymbol{J}}^{M}_{t}=-[{\boldsymbol{J}}^{M_{1}}_{t}\;{\boldsymbol{J}}^{M_{2}}_{t}\;{\boldsymbol{J}}^{M_{3}}_{t}]$, where ${\boldsymbol{J}}^{M_{1}}_{t}={\boldsymbol{r}}_{t}^{M}$, ${\boldsymbol{J}}^{M_{3}}_{t}=[l_{c3}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}\;l_{c3}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}]^{T}$, and
@@
In the above model for quiet stance, $l_{ci}=\frac{\frac{1}{2}m_{i}l_{i}+\sum_{j=i+1}^{3}m_{i}l_{j}}{m^{t}_{t}}$, $i=1,2,3$, and $m^{t}_{t}=m_{1}+m_{2}+m_{3}$. For kneeling gait, $l_{ci}=\frac{\frac{1}{2}m_{i}l_{i}+\sum_{j=i+1}^{3}m_{i}l_{j}}{m^{t}_{d}}$, $i=2,3$, $m^{t}_{d}=m_{2}+m_{3}$ The acceleration of the COM is obtained as ${\boldsymbol{\ddot{r}}}^{M}_{i}=[{\boldsymbol{\dot{J}}}^{M}_{i}\;{\boldsymbol{J}}^{M}_{i}]\dot{{\boldsymbol{x}}}_{i}=\mathcal{J}_{i}^{M}\dot{{\boldsymbol{x}}}_{i}$, where $\mathcal{J}_{i}^{M}=[{\boldsymbol{\dot{J}}}^{M}_{i}\;{\boldsymbol{J}}^{M}_{i}]$, $i=t,d$. The GRFs in the horizontal and vertical directions are respectively computed as
$F^{x}_{i}=m^{t}_{i}\ddot{x}^{m}_{i},\quad F^{z}_{i}=m^{t}_{i}(\ddot{z}^{m}_{i}+g),\,i=t,d,$ (3)

### v2

@@ Abstract
Construction workers experience serious safety and health risks in hazardous working environments. Quiet stance and kneeling are among the most common postures performed by construction workers during their daily work. This paper first analyzes lower-limb joint influence on neural balance control strategies using the frequency behavior of the intersection point of ground reaction forces. To evaluate the impact of elevation and wearable knee exoskeletons on postural balance and welding task performance, we design virtual- and mixed-reality (VR/MR) to simulate elevated environments and welding tasks. A linear quadratic regulator-controlled triple- and double-link inverted pendulum model is used for balance strategy quantification in stance and kneeling, respectively. Extensive multi-subject experiments are conducted to evaluate the usability of wearable exoskeletons in destabilizing construction environments. The quantified balance strategies capture the significance of the knee joint during balance control of quiet stance and kneeling gaits. Results show that center of pressure sway area is reduced up to 62% in quiet stance and 39% in kneeling gait for subjects tested in high-elevation worksites with knee exoskeleton assistance. The balance and multitask evaluation confirm and provide guidance on exoskeleton design to mitigate the fall risk in construction.
@@
Construction workers commonly perform tasks that require prolonged quiet stance or kneeling gaits on high elevations. Worker balance can be undermined by chronic knee injuries, musculoskeletal disorders, and destabilizing visual perturbations caused by occupational activities. Wearable knee exoskeletons have evolved as promising interventions to reduce knee joint stress across a variety of work gaits in construction. Emerging technologies such as virtual- and mixed-reality (VR/MR) provide an enabling tool to study underlying balance strategies to complete tasks in dynamic environments. The VR/MR- generated immersive elevated welding environment are leveraged to examine the effects of threatening visual stimuli, wearable exoskeletons, and construction tasks on worker balance and skill performance. Intersection point height frequency analysis is used to quantify the neural balance strategy during various testing scenarios. We particularly explore the often-neglected role of the knee joint to facilitate research on knee exoskeleton assisted balance in stance and kneeling gaits. The experimental results provide insight into the efficacy of knee exoskeletons in improving worker’s stability and task performance in construction. The results highlight the need for a holistic approach to exoskeleton design to ensure that developed solutions can be safely and successfully integrated into the workplace environment.
@@ I Introduction
Construction workers commonly perform tasks that require prolonged quiet stance or kneeling gaits in high elevations or cluttered environments. These awkward gaits in construction increase the risk of work-related musculoskeletal disorders (WMSDs), such as knee pain or knee osteoarthritis [1]. In both stance and kneeling gaits, human workers are inherently unstable in an upright posture, requiring continuous engagement of muscles to regulate joint torques and maintain postural stability [2, 3]. Although the neural balance mechanism at quiet stance has been extensively studied, the underlying control strategy remains elusive [4]. Previous studies have investigated the ankle or hip strategies during quiet standing [5, 6], and most of them used the single-link inverted pendulum model to analyze human balance strategies. The importance of the knee and hip joints at quiet stance was demonstrated in [7, 8]. The knee joint contribution to maintaining an upright stance is however often overlooked. Few research work was reported for postural control of kneeling gaits. In [9], a comparative analysis of quiet kneeling and stance postures was conducted with a focus on the role of visual feedback during kneeling. A single inverted pendulum model was used in [6] to study the influences of elevation on the postural balance of kneeling gait.
@@
All of the works above primarily use the motion of the center of mass (COM) or center of pressure (COP) of the foot-ground contact to quantify the human balance performance. Metrics such as power spectral density (PSD) of the COP and acceleration were used to compare and discriminate between quiet stance and kneeling [9, 6]. The metrics solely derived from COP or COM are limited in capturing the balance dynamics and control as they are unable to account for both translational and rotational body accelerations simultaneously. To address these limitations, the study in [10, 4] introduced the intersection point (IP) of the ground reaction forces (GRF) during quiet stance. In [11], a double-link inverted pendulum (DIP) model was used to quantify the frequency-dependence of the IP height during quiet stance. The study focused on quantifying ankle and hip joint strategies and emphasized the importance of multiple joint-level contributions to overall balance control.
@@
Among the diverse environmental factors affecting balance control for construction workers, elevation is notably significant [12, 13, 14]. Acrophobia exacerbates this challenge, as it leads individuals to overestimate heights, intensifying fear and triggering physiological reactions that undermine balance [15]. Additionally, visual and environmental perturbations further compromise postural stability [16]. To address these challenges, wearable devices like knee exoskeletons have emerged as pivotal fall risk interventions [17, 18, 19, 20]. However, few work has been reported on knee-based balance augmentation for stance and kneeling. Virtual- and mixed-reality (VR/MR) has been used to simulate high-elevation conditions and study neural balance control among construction workers [21, 6]. The work in [22] showed safety reinforcement training on fall prevention for construction workers using VR scenes. The work in [23] conducted neurophysiological assessments for welding task performance at virtual height to study cognitive effects on balance. A comprehensive review of VR/MR applications in construction safety can be found in [24].
@@
One main goal of this work is to study the influence of lower-limb joints, particularly the often-overlooked knee joint, on neural balance control in stance and kneeling gaits. We also evaluate the impact of elevation and wearable knee exoskeletons on postural balance and welding task performance in a realistic construction environment. Similar to previous studies (e.g., [25, 26, 11]), a linear quadratic regulator (LQR) is used as the human neural balance controller. Triple- and double-link inverted pendulum (TIP and DIP) models are taken to capture the joint-level effects for quiet stance and kneeling gaits, respectively. The models are used to quantify the observed neural balance control strategy from the frequency characteristics of the IP heights. The LQR allows for the explicit consideration of different parameters on balance performance, facilitating the interpretation of results in terms of physiological and biomechanical implications [11]. Human subject experiments are conducted using incorporating VR/MR environments that simulate realistic welding worksites at varying elevations to evaluate the influence of elevation on postural balance and task performance. The subjects are outfitted with a wearable knee exoskeleton to test the usability and functionality of the intervention. The combination of experimental data and model analysis helps the extraction of physiologically plausible insights into the balance strategy chosen by subjects. Both objective metrics and human subjective feedback confirm that the exoskeletons have promising potentials to improve both balance and task execution efficiency in construction.
@@
The main contributions of this work are twofold. First, the study introduces new modeling and analysis for neural balance control that accounts for knee joint strategy with the IP height as the performance metric for both quiet stance and kneeling. To the authors’ best knowledge, no previously reported work elucidates the involvement of knee joint strategy in stance or kneeling gaits. Second, this study provides a comprehensive evaluation of how elevation and wearable exoskeletons affect balance and task performance in construction. The quantitative findings offer valuable insights for the development of balance-enhanced wearable robotic interventions. This study integrates VR/MR to simulate realistic construction trades, which improves the accuracy of assessing exoskeleton-assisted balance and skilled task performance. Compared with the previous conference presentation [27] that focused only on quiet stance, this work extends to kneeling gaits by introducing the TIP/DIP models to investigate the joint-level effects on postural balance control. The use of the LQR neural controller allows us to analyze the hip, knee, and ankle joints’ contributions specifically to stance and kneeling balance. The analysis and comparison with experiments are new and contribute to a broader understanding of different work postures in construction.
@@
The rest of the paper is organized as follows. We introduce the multi-link models and neural balance control in Section II. The experimental setup and protocols are presented in Section III. Data analysis and processing pipelines are explained in Section IV. Section V presents the experimental results, followed by discussions in Section VI. Finally, we give the concluding summary in Section VII.
@@ II Multi-Link Models and Balance Control
In this section, we first present the inverted pendulum models that describe quiet stance and kneeling, respectively. Next, the IP for human postural balance is discussed and we then present the LQR-based neural controller.
@@ II-A Biomechanical Models for Stance and Kneeling
Fig. 1 illustrates the welder’s stance posture and schematic consisting of three interconnected segments. The upper segment, which includes the head, arms, and trunk, is connected to the middle segment, representing the thigh, and the lower segment, representing the shank. Fig. 1 shows the welder’s working pose when kneeling on the ground with both legs alongside the schematic of the DIP model. In the case of quiet stance, the ankle joint is assumed to be the pin joint on the ground, whereas for kneeling, the knee joint is assumed to be the pin joint on the ground and feet are assumed to support no weights. Several assumptions are considered to model quiet stance and kneeling gaits. Small angle approximation is used by assuming that body movements remain small. We focus on planar motion in only the sagittal plane. The human body is modeled as a series of rigid segments (via DIP and TIP models). Additionally, we assume that human workers maintain a stationary posture around the unstable equilibrium with only the foot or knee in contact with the ground, in other words, we do not consider multi-contact postures such as when workers lean forward or use their hands for additional support.
The joint angles and joint torques at the ankle, knee, and hip are denoted as $\theta_{i}$ and $\tau_{i}$, $i=1,2,3$, respectively. The masses and mass moments of inertia about the centers of each link are denoted as $m_{i}$ and $I_{i}$, $i=1,2,3$, respectively. The distance between the mass center and the lower end of each link is denoted as $l_{c_{i}}$, $i=1,2,3$, respectively. Defining ${\boldsymbol{\theta}}_{t}=[\theta_{1}\;\theta_{2}\;\theta_{3}]^{T}$ and ${\boldsymbol{\theta}}_{d}=[\theta_{2}\;\theta_{3}]^{T}$, the equations of motion for the TIP/DIP models are written as
@@
Introducing ${\boldsymbol{x}}_{i}=[{\boldsymbol{\theta}}_{i}^{T}\;{\boldsymbol{\dot{\theta}}}_{i}^{T}]^{T}$, $i=t,d$, from (1), we obtain
@@
The position vectors of the COM are denoted as ${\boldsymbol{r}}^{M}_{i}=[x^{m}_{i}\,z^{m}_{i}]^{T}$, $i=t,d$, for stance and kneeling gaits, respectively. For stance posture, $x^{m}_{t}=-l_{c1}\mathop{\rm s}\nolimits_{\theta_{1}}-l_{c2}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}}-l_{c3}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}$ and $z^{m}_{t}=l_{c1}\mathop{\rm c}\nolimits_{\theta_{1}}+l_{c2}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}}+l_{c3}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}$, where notations $\mathop{\rm s}\nolimits_{\theta_{i}}=\sin\theta_{i}$, $\mathop{\rm c}\nolimits_{\theta_{i}}=\cos\theta_{i}$, $\mathop{\rm s}\nolimits_{\theta_{i}+\theta_{j}}=\sin(\theta_{i}+\theta_{j})$, and $\mathop{\rm c}\nolimits_{\theta_{i}+\theta_{j}}=\cos(\theta_{i}+\theta_{j})$ for angles $\theta_{i}$ and $\theta_{j}$, $i\neq j$, $i,j=1,2,3$. For quiet stance, we define the Jacobian matrix of the COM to ${\boldsymbol{x}}_{i}$ as ${\boldsymbol{J}}^{M}_{t}=-[{\boldsymbol{J}}^{M_{1}}_{t}\;{\boldsymbol{J}}^{M_{2}}_{t}\;{\boldsymbol{J}}^{M_{3}}_{t}]$, where ${\boldsymbol{J}}^{M_{1}}_{t}={\boldsymbol{r}}_{t}^{M}$, ${\boldsymbol{J}}^{M_{3}}_{t}=[l_{c3}\mathop{\rm c}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}\;l_{c3}\mathop{\rm s}\nolimits_{\theta_{1}+\theta_{2}+\theta_{3}}]^{T}$, and
@@
In the above model for quiet stance, $l_{ci}=\frac{\frac{1}{2}m_{i}l_{i}+\sum_{j=i+1}^{3}m_{i}l_{j}}{m^{t}_{t}}$, $i=1,2,3$, and $m^{t}_{t}=m_{1}+m_{2}+m_{3}$. For kneeling gait, $l_{ci}=\frac{\frac{1}{2}m_{i}l_{i}+\sum_{j=i+1}^{3}m_{i}l_{j}}{m^{t}_{d}}$, $i=2,3$, $m^{t}_{d}=m_{2}+m_{3}$ The acceleration of the COM is obtained as ${\boldsymbol{\ddot{r}}}^{M}_{i}=[{\boldsymbol{\dot{J}}}^{M}_{i}\;{\boldsymbol{J}}^{M}_{i}]\dot{{\boldsymbol{x}}}_{i}=\mathcal{J}_{i}^{M}\dot{{\boldsymbol{x}}}_{i}$, where $\mathcal{J}_{i}^{M}=[{\boldsymbol{\dot{J}}}^{M}_{i}\;{\boldsymbol{J}}^{M}_{i}]$, $i=t,d$. The GRFs in the horizontal and vertical directions are computed as
$F_{x}^{i}=m^{t}_{i}\ddot{x}^{m}_{i},\quad F_{z}^{i}=m^{t}_{i}(\ddot{z}^{m}_{i}+g),\,\,i=t,d,$ (3)